# Human Motion Capture from Loose and Sparse Inertial Sensors with Garment-aware Diffusion Models

### v1

@@ Abstract
Motion capture using sparse inertial sensors has shown great promise due to its portability and lack of occlusion issues compared to camera-based tracking. Existing approaches typically assume that IMU sensors are tightly attached to the human body. However, this assumption often does not hold in real-world scenarios. In this paper, we present a new task of full-body human pose estimation using sparse, loosely attached IMU sensors. To solve this task, we simulate IMU recordings from an existing garment-aware human motion dataset. We developed transformer-based diffusion models to synthesize loose IMU data and estimate human poses based on this challenging loose IMU data. In addition, we show that incorporating garment-related parameters while training the model on simulated loose data effectively maintains expressiveness and enhances the ability to capture variations introduced by looser or tighter garments. Experiments show that our proposed diffusion methods trained on simulated and synthetic data outperformed the state-of-the-art methods quantitatively and qualitatively, opening up a promising direction for future research.
@@ 1 Introduction
Figure 1: We present a novel task of full-body human pose estimation using sparse IMUs that are loosely attached to garments. To address this challenge, we propose methods for simulating and synthesizing loose IMU signals using diffusion models. Leveraging both simulated and synthetic data, we design diffusion models conditioned on IMU data alone or IMU data and clothing model parameters, enabling robust human motion estimation. We evaluate the effectiveness of our approach across three datasets: a simulated dataset, a real-world upper-body dataset, and a real-world full-body dataset.
@@
    a transformer-based diffusion model that estimates full-body poses from loosely attached IMUs. We show that incorporating garment-related parameters significantly enhances pose accuracy, even amid variations in signal noise and differences in garment fit across individuals.
@@ 2 Related Work
To the best of our knowledge, the only method tailored for loosely worn IMUs is Loose Inertial Poser (LIP) Zuo et al. (2024). It introduces a Secondary Motion Autoencoder that treats secondary motion as Gaussian noise, using latent space learning to model variation across body shapes and garment types. A temporal coherence mechanism ensures smooth, realistic motion. Trained on AMASS and evaluated on real upper-body data, LIP achieved $<20$° joint rotation error and 10.6 cm position error and outperformed several SOTA models (PIP, TIP, DIP, TransPose) trained on simulated loose IMU data. LIP uses LSTMs, which suit sequential IMU inputs. In contrast, our pipeline leverages modern generative models to better capture uncertainty and complex motion patterns.
@@
3 Proposed Method
@@ 3.1 Method Overview
Training data is generated by simulating IMU signals from a garment-aware motion capture dataset. Our approach uses two conditional diffusion models: a primary model for pose estimation and a secondary model for generating realistic loose-wear IMU signals. The secondary model, conditioned on tight-wear IMU data and pose, synthesizes loose-wear signals. These are used to train the primary model, which predicts pose, tight-wear IMU signals, and joint positions. A garment-aware variant further conditions on clothing parameters to enhance realism.
@@ 3.4 Conditional Diffusion Models for Pose Estimation
To estimate the pose based on the observations from loosely attached sensors, we developed a conditional diffusion model (Figure 4). These sensor observations serve as conditions throughout the diffusion process, enabling the model to generate not only pose estimates but also predictions of tight IMU data and joint positions. Given the conditional nature of this model, where predictions fully rely on the provided input, a flexible number of available sensors cannot be accommodated. Therefore, we propose two distinct models—one for the upper-body pose estimation and another for the full-body pose estimation.
The upper-body model predicts the positions and rotations of 14 upper-body joints, including leaf joints, while the whole-body model predicts all 24 joints, as defined by the SMPL body model Loper et al. (2015). The loss function for the upper-body model is defined as
@@
The loss function for the whole-body model follows the same structure, with $\mathcal{L}_{p_{al}}$ (position error of arms and legs) replacing $\mathcal{L}_{p_{a}}$ (position error of arms) to account for the position error of all extremities. Similarly, $\mathcal{L}_{r_{j}}$ now incorporates the rotation estimation errors for the lower body as well. All loss functions are based on L1 loss. Additionally, we assign higher weights to the root joint’s rotation and the positions of the extremities, as these proved to be particularly challenging for accurate estimation.
@@ 3.5 Garment-Aware Conditional Model
The conditional model trained on real-world whole-body IMU data revealed issues with very loose garment configurations, particularly with the root joint sensor, which is placed on top of the jacket and over the pants, leading to highly noisy recordings. Similarly, sensors placed on the extremities often provide misleading observations due to garment movement. To mitigate this intrinsic uncertainty, we try to enhance the conditional model by integrating garment-related information. To achieve this, we utilize seven simulations generated using TailorNet’s MaleShirt model. Six of these simulations correspond to a TallThin physique, denoted as [180, 22], where 180 represents height in centimeters and 22 refers to the body mass index (BMI). These simulations encompass six garment style variations, with $\gamma$ values of 0, 5, 10, 15, 20, and 24. The final simulation represents a ShortFat physique, denoted as [160, 30], with a $\gamma$ value of 0. We focus on four upper-body IMU sensors for these simulations.
@@ 4.3 Evaluation Results
In this section, we compare our proposed methods and the state-of-the-art pose estimation methods using loosely attached IMUs. Note that LIP has previously demonstrated superior performance over methods designed for tightly-attached IMUs Mollyn et al. (2023); Yi et al. (2022); Huang et al. (2018); Yi et al. (2021); Jiang et al. (2022b).
@@
Figures 5, 6 and 7 illustrate the qualitative performance of our conditional models, including the upper-body and whole-body garment-unaware models, as well as the upper-body garment-aware model.
@@
Based on the results presented in Table 2, we observe that models trained on simulated and synthetic data perform exceptionally well on the five AMASS datasets. However, their performance gradually declines on the LIP dataset, culminating in significantly reduced accuracy on the whole-body real-world dataset. In contrast, models trained on real-world data achieve good performance on the whole-body dataset but struggle with the AMASS and LIP datasets, likely due to the tighter fit conditions. This indicates that all diffusion models and both LIP models perform significantly better when the distribution of test data closely matches that of the training data, revealing a problem with adaptation to different garment conditions or body shapes.
In very challenging scenarios with noisy IMU observations, the models trained on real-world data (Figures 5 and 6) tend to converge towards the mean predictions (neutral pose with legs and arms in relaxed positions). This raises the question of whether the data augmentation approach and emphasis on realistic data are optimal, given that noise and ambiguity in the recorded signals not only mislead the model but also reduce motion dynamics.
@@
While we introduced height and BMI to determine whether a TallThin or ShortFat physique was used for the clothing simulation, it remains to be seen whether these parameters in real-world scenarios pertain more to garment-related characteristics or the participant’s body shape. Since we lacked detailed information about the participants’ physiques, we were unable to assess this aspect or evaluate each participant’s performance across a long sequence and varying garment styles. However, our experiments on shorter motion sequences, using the optimal selection of the three parameters, demonstrate an improved quantitative performance, as well as a clear qualitative advantage. This is illustrated in Figure 7. Both suitable and unsuitable garment-related information can impact pose estimation accuracy, potentially improving or degrading it, with joint angle errors fluctuating by ±5 degrees and joint position errors by up to ±10 cm (compared to the garment-unaware model). Finally, the selection of the three parameters began by choosing values within the training range—height between 160 and 180 cm, BMI between 22 and 30, and the $\gamma$ parameter taking one of six predefined values. When the parameters were assigned unrealistic real-world height and BMI values, the motion became more restricted and less dynamic, resembling the behavior of the model trained on real-world data with loosely fitting garments (Upper-body Real).
@@ 4.4 Ablation Study
We attribute our model’s generalization ability and reduced jitter to the inclusion of this loss function. It compels the model to generate similar outputs for slightly varied and noisier versions of the observations, thereby enhancing its robustness. This is demonstrated by the degraded performance on real-world datasets and the enhanced performance on simulated datasets—which share the same characteristics as the training dataset—when the loss is omitted from the objective function. Table 4 presents the performance of the upper- and whole-body models trained on simulated data without this loss function.
@@
Real-world upper-body dataset MPJRE 21.701 ± 6.390 __
MPJPE 13.214 ± 7.645 __
MPJVE 43.738 __
Jitter 289.034 __
GT Jitter 33.412 __
@@
Table 4: Performance metrics of the conditional models trained on simulated data without $\mathcal{L}_{\textit{c}}$.
@@
The objective function mirrors that of the conditional whole-body model, with the addition of the L1 loss term, $\mathcal{L}_{\text{loose\_recon}}$, to account for errors in the reconstruction of the missing root sensor’s data. While including all observations led to similar performance in both the conditional and unconditional models, the results in Table 5 suggest that completely excluding the root sensor’s data from all frames during evaluation is not advisable, as this data is not consistently unreliable. The presence of reliable observations is likely a contributing factor to the better performance of the model when all data is available. By omitting certain data entirely, the model is granted the freedom to predict the pose of the affected joint by considering other sensors’ data that could be misleading itself (such as data obtained from the extremities), which can lead to inaccurate estimates, particularly pronounced when root rotation is minimal.
@@
Table 5: Comparison between the conditional whole-body model (trained/evaluated on all six sensors) and an inpainting-based unconditional diffusion model (trained/evaluated with missing root sensor).
@@ 5 Conclusion
In this paper, we introduce a new task of full-body pose estimation from sparse and loosely attached IMU sensor data. Our approach builds on a multi-source training strategy that leverages simulated, synthetic, and real-world IMU data. We simulate IMU observations using garment-aware motion datasets and develop a diffusion model to synthesize realistic loose-wear IMU signals. This framework establishes a foundation for a comprehensive comparison of model performance across different data sources, allowing us to systematically evaluate the strengths and limitations of each. In addition, we demonstrate that incorporating garment-related parameters during training on simulated loose data enhances model expressiveness and helps address the inherent ambiguities of loosely attached IMU sensors. Experiments show our diffusion models outperform the state-of-the-art methods, providing a new direction for future research.

### v2

@@ Abstract
Motion capture using sparse inertial sensors has shown great promise due to its portability and lack of occlusion issues compared to camera-based tracking. Existing approaches typically assume that IMU sensors are tightly attached to the human body. However, this assumption often does not hold in real-world scenarios. In this paper, we present Garment Inertial Poser (GaIP), a method for estimating full-body poses from sparse and loosely attached IMU sensors. We first simulate IMU recordings using an existing garment-aware human motion dataset. Our transformer-based diffusion models synthesize loose IMU data and estimate human poses from this challenging loose IMU data. We also demonstrate that incorporating garment-related parameters during training on loose IMU data effectively maintains expressiveness and enhances the ability to capture variations introduced by looser or tighter garments. Our experiments show that our diffusion methods trained on simulated and synthetic data outperform state-of-the-art inertial full-body pose estimators, both quantitatively and qualitatively, opening up a promising direction for future research on motion capture from such realistic sensor placements.
@@ 1 Introduction
Figure 1: Garment Inertial Poser (GaIP) is a novel model for estimating full-body poses from sparse inertial sensors that are loosely attached to garments. While sensors that are attached to clothing rather than tightly strapped to the body is much more realistic, it is simultaneously more challenging. We first introduce diffusion models for simulating and synthesizing loose IMU signals. Leveraging this data, we propose diffusion models conditioned on IMU data alone or IMU data and clothing model parameters, enabling robust human motion estimation. We evaluate the effectiveness of our approach across three datasets: a simulated dataset, a real-world upper-body dataset, and a real-world full-body dataset.
@@
    a transformer-based diffusion model Garment Inertial Poser (GaIP) that estimates full-body poses from loosely attached IMUs. We show that incorporating garment-related parameters significantly enhances pose accuracy, even amid variations in signal noise and differences in garment fit across individuals.
@@ 2 Related Work
To the best of our knowledge, the only method tailored for loosely worn IMUs is Loose Inertial Poser (LIP) Zuo et al. (2024). It introduces a Secondary Motion Autoencoder that treats secondary motion as Gaussian noise, using latent space learning to model variation across body shapes and garment types. A temporal coherence mechanism ensures smooth, realistic motion. Trained on AMASS and evaluated on real upper-body data, LIP achieved <20<20° joint rotation error and 10.6 cm position error and outperformed several SOTA models (PIP, TIP, DIP, TransPose) trained on simulated loose IMU data. LIP uses LSTMs, which suit sequential IMU inputs. In contrast, our pipeline leverages modern generative models to better capture uncertainty and complex motion patterns.
@@
3 Garment Inertial Poser (GaIP) Method
@@ 3.1 Method Overview
We generate training data by simulating IMU signals from a garment-aware motion capture dataset. Our approach uses two conditional diffusion models: a primary model for pose estimation and a secondary model for generating realistic loose-wear IMU signals. The secondary model, conditioned on tight-wear IMU data and pose, synthesizes loose-wear signals to train our primary GaIP model. It predicts pose, tight-wear IMU signals, and joint positions. A garment-aware variant further conditions on clothing parameters to enhance realism.
@@ 3.4 Conditional Diffusion Models for Pose Estimation
To estimate the pose based on the observations from loosely attached sensors, we developed GaIP, a conditional diffusion model (Figure 4). These sensor observations serve as conditions throughout the diffusion process, enabling the model to generate not only pose estimates but also predictions of tight IMU data and joint positions. Given the conditional nature of this model, where predictions fully rely on the provided input, a flexible number of available sensors cannot be accommodated. Therefore, we propose two distinct models—one for the upper-body pose estimation (GaIP Upper-body) and another for the full-body pose estimation (GaIP Full-body).
GaIP Upper-body predicts the positions and rotations of 14 upper-body joints, including leaf joints, while GaIP Full-body predicts all 24 joints, as defined by the SMPL body model Loper et al. (2015). The loss function for the upper-body model is defined as
@@
The loss function for the whole-body model GaIP Full-body follows the same structure, with $\mathcal{L}_{p_{al}}$ (position error of arms and legs) replacing $\mathcal{L}_{p_{a}}$ (position error of arms) to account for the position error of all extremities. Similarly, $\mathcal{L}_{r_{j}}$ now incorporates the rotation estimation errors for the lower body as well. All loss functions are based on L1 loss. Additionally, we assign higher weights to the root joint’s rotation and the positions of the extremities, as these proved to be particularly challenging for accurate estimation.
@@ 3.5 Garment-Aware Conditional Model
The conditional model trained on real-world whole-body IMU data revealed issues with very loose garment configurations, particularly with the root joint sensor, which is placed on top of the jacket and over the pants, leading to highly noisy recordings. Similarly, sensors placed on the extremities often provide misleading observations due to garment movement. To mitigate this intrinsic uncertainty, we try to enhance GaIP by integrating garment-related information. To achieve this, we utilize seven simulations generated using TailorNet’s MaleShirt model. Six of these simulations correspond to a TallThin physique, denoted as [180, 22], where 180 represents height in centimeters and 22 refers to the body mass index (BMI). These simulations encompass six garment style variations, with $\gamma$ values of 0, 5, 10, 15, 20, and 24. The final simulation represents a ShortFat physique, denoted as [160, 30], with a $\gamma$ value of 0. We focus on four upper-body IMU sensors for these simulations.
@@ 4.3 Evaluation Results
In this section, we compare our GaIP models and the state-of-the-art pose estimation methods using loosely attached IMUs. Note that LIP has previously demonstrated superior performance over methods designed for tightly-attached IMUs Mollyn et al. (2023); Yi et al. (2022); Huang et al. (2018); Yi et al. (2021); Jiang et al. (2022b).
@@
Figures 5, 6 and 7 illustrate the qualitative performance of our GaIP models, including the upper-body and whole-body garment-unaware models, as well as the upper-body garment-aware model.
@@
Based on the results presented in Table 2, we observe that GaIP models trained on simulated and synthetic data perform exceptionally well on the five AMASS datasets. However, their performance gradually declines on the LIP dataset, culminating in significantly reduced accuracy on the whole-body real-world dataset. In contrast, GaIP models trained on real-world data achieve good performance on the whole-body dataset but struggle with the AMASS and LIP datasets, likely due to the tighter fit conditions. This indicates that all GaIP diffusion models and both LIP models perform significantly better when the distribution of test data closely matches that of the training data, revealing a problem with adaptation to different garment conditions or body shapes.
In very challenging scenarios with noisy IMU observations, GaIP models trained on real-world data (Figures 5 and 6) tend to converge towards the mean predictions (neutral pose with legs and arms in relaxed positions). This raises the question of whether the data augmentation approach and emphasis on realistic data are optimal, given that noise and ambiguity in the recorded signals not only mislead the model but also reduce motion dynamics.
@@
While we introduced height and BMI to determine whether a TallThin or ShortFat physique was used for the clothing simulation, it remains to be seen whether these parameters in real-world scenarios pertain more to garment-related characteristics or the participant’s body shape. Since we lacked detailed information about the participants’ physiques, we were unable to assess this aspect or evaluate each participant’s performance across a long sequence and varying garment styles. However, our experiments on shorter motion sequences, using the optimal selection of the three parameters, demonstrate an improved quantitative performance, as well as a clear qualitative advantage. This is illustrated in Figure 7. Both suitable and unsuitable garment-related information can impact pose estimation accuracy, potentially improving or degrading it, with joint angle errors fluctuating by ±5 degrees and joint position errors by up to ±10 cm (compared to the garment-unaware model). Finally, the selection of the three parameters began by choosing values within the training range—height between 160 and 180 cm, BMI between 22 and 30, and the $\gamma$ parameter taking one of six predefined values. When the parameters were assigned unrealistic real-world height and BMI values, the motion became more restricted and less dynamic, resembling the behavior of the model trained on real-world data with loosely fitting garments (GaIP Upper-body Real).
@@ 4.4 Ablation Study
We attribute GaIP’s generalization ability and reduced jitter to the inclusion of this loss function. It compels the model to generate similar outputs for slightly varied and noisier versions of the observations, thereby enhancing its robustness. This is demonstrated by the degraded performance on real-world datasets and the enhanced performance on simulated datasets—which share the same characteristics as the training dataset—when the loss is omitted from the objective function. Table 4 presents the performance of the upper- and whole-body models trained on simulated data without this loss function.
@@
Real-world upper-body dataset MPJRE 21.701 ± 6.390 $_$
MPJPE 13.214 ± 7.645 $_$
MPJVE 43.738 $_$
Jitter 289.034 $_$
GT Jitter 33.412 $_$
@@
Table 4: Performance metrics of GaIP models trained on simulated data without $\mathcal{L}_{\textit{c}}$.
@@
The objective function mirrors that of GaIP Full-body model, with the addition of the L1 loss term, $\mathcal{L}_{\text{loose\_recon}}$, to account for errors in the reconstruction of the missing root sensor’s data. While including all observations led to similar performance in both the conditional and unconditional models, the results in Table 5 suggest that completely excluding the root sensor’s data from all frames during evaluation is not advisable, as this data is not consistently unreliable. The presence of reliable observations is likely a contributing factor to the better performance of the model when all data is available. By omitting certain data entirely, the model is granted the freedom to predict the pose of the affected joint by considering other sensors’ data that could be misleading itself (such as data obtained from the extremities), which can lead to inaccurate estimates, particularly pronounced when root rotation is minimal.
@@
Table 5: Comparison between GaIP Full-body model (trained/evaluated on all six sensors) and an inpainting-based unconditional diffusion model (trained/evaluated with missing root sensor).
@@ 5 Conclusion
In this paper, we introduce Garment Inertial Poser—a novel method that estimates full-body pose from sparse and loosely attached IMU sensor data. Our approach builds on a multi-source training strategy that leverages simulated, synthetic, and real-world IMU data. We simulate IMU observations using garment-aware motion datasets and develop a diffusion model to synthesize realistic loose-wear IMU signals. This framework establishes a foundation for a comprehensive comparison of model performance across different data sources, allowing us to systematically evaluate the strengths and limitations of each. In addition, we demonstrate that incorporating garment-related parameters during training on simulated loose data enhances model expressiveness and helps address the inherent ambiguities of loosely attached IMU sensors. Experiments show GaIP outperforms the state-of-the-art methods, providing a new direction for future research.